Two angles are complementary. The measure of one angle exceeds two times the measure of the other angle by 21. What is the degree of the smaller angle? Could you please show the work on how to solve this. Thanks!
Found 2 solutions by josgarithmetic, lwsshak3:
Answer by josgarithmetic(39625)

Measures x and y.
x+y=90 because complementary;
y-2x=21 through the description;
Solve.

Answer by lwsshak3(11628)

Two angles are complementary. The measure of one angle exceeds two times the measure of the other angle by 21. What is the degree of the smaller angle? Could you please show the work on how to solve this.
***
let x=larger angle
let y=smaller angle
x+y=90
y=90-x
..
x=2y+21
x=2(90-x)+21
x=180-2x+21
3x=201
x=67˚
y=90-x=23˚
smaller angle=23˚